hello, I have a story problem I could use some help with if you have time, A bag contains 6 red marbles, 9 blue marbles, and 5 green marbles. You withdrew one marble, replaced it, and then withdrew another marble. What is the probabilty that you do NOT pick 2 green marbles?
Answer by venugopalramana(3286)

A bag contains 6 red marbles, 9 blue marbles, and 5 green marbles. You withdrew one marble, replaced it, and then withdrew another marble. What is the probabilty that you do NOT pick 2 green marbles?
IN SUCH PROBLEMS IT IS EASIER TO FIND PROBABILITY OF DRAWING 2 GREEN MARBLES AND SUBTRACT IT FROM 1 TO GET THE PROBABILITY OF NOT DRAWING 2 GREEN MARBLES.
IN THE 1ST.TRIAL PROBABILITY OF DRAWING GREEN MARBLE =5/(6+9+5)=5/20=1/4
SINCE THE MARBLE IS REPLACED AFTER 1ST.WITHDRAWAL THE PROBABILITY OF DERAWING GREEN MARBLE AT SECOND ATTEMPT IS ALSO SAME AS FIRST =1/4
HENCE PROBABILITY OF DRAWING 2 GREENS =(1/4)*(1/4)=1/16
HENCE OF PROBABILITY OF NOT DRAWING 2 GREEN MARBLES =1-1/16=15/16